# Detecting High-Stakes Interactions with Activation Probes

### v2

@@ A.2 Baselines
For our baselines, we designed four different prompt templates and evaluated the baselines using all four templates on the development split of our evaluation datasets. For each model, we then picked the template that led to the highest average AUROC (see Figure 10) and evaluated it on the test split of the evaluation datasets. Note that, especially for smaller models, the choice of prompt template can have a great effect on the performance of prompted baselines.
The prompt templates we chose for each model can be found in Table 4. Below you can find the four prompt templates:
@@
Llama-1b default
Llama-8b default
Llama-70b single-word
Gemma-1b prompt-at-end
Gemma-12b single-word
Gemma-27b single-word
@@
Hyperparameters used for training our finetuned baselines can be found in Figure 8. We computed a validation loss on the validation split of our synthetic dataset to decide which checkpoint to use. Note that we while we trained for 10-25 epochs for most models, the best epoch (and checkpoint we used for evaluation) was between 2-4 for smaller models (Llama-1b and Gemma-1b), 3-5 for Llama-8b and 1-2 for Gemma-12b. In particular, learning rate scheduling did not help for smaller models. For Llama-8b, we found gradient clipping to be particularly important for getting good results. Surprisingly, for Gemma-12b, a gradient accumulation of 4 led to much better performance on the dev split than using a gradient accumulation of 128, mostly because of the ToolACE dataset (AUROC of around 0.9 compared to AUROC below 0.7).
@@ A.3 Estimating Computational Cost
  • Activations from Llama-70b for the given conversation (in exactly that conversation format) are already available at no extra cost. This is because we assume the monitored system already uses Llama-70b as its primary model to generate responses to user queries, and the activations we need are a natural byproduct of this inference process.
    Note that the prompted baseline of Llama-70b uses a different prompt structure, therefore we assume that this baseline model has to process the whole sequence. (The cost of the prompted Llama-70b could potentially be reduced by keeping the same conversation format and asking about the stakes label at the end, but it is unclear how much that would affect performance.)
@@
Plugging in the parameters for the specific models, we obtain the compute costs formulas listed in Table 5.
Model $N$ $m$ $d$ $w$ Compute Cost Formula
@@
Note that finetuned and prompted baselines use different prompt templates, which explains why the number of FLOPs for the finetuned and prompted baseline of the same model can be different (e.g. see Figure 13).
@@
Note that we estimate the number of FLOPs for our probes as well, but they are several orders of magnitude smaller and thus not noticeable in probe+baseline combinations in the plot (Figure 6).
The specific estimates for the relevant probes can be found in Table 6.
@@
Table 6: Compute cost formulas for different probes on Llama-70b activations as functions of input sequence length $S$. Note that activations of Llama-70B have size $d=8192$.
@@ A.4 Hardware Resources
  • Cloud servers from the cloud provider Lambda Labs, Vast.ai and Runpod were used for our experiments. For computing activations, a machine with 2xH100 GPUs (80GB each) and $\sim 400$GB of RAM was used. For training and evaluating probes with pre-computed activations, a machine with 1xA100 (40GB) and $\sim 200$GB of RAM was sufficient. For finetuning baselines up to 8B parameters, a machine with 1xH100 (80GB) was used. Gemma-12B was finetuned on a machine with 1xH100 NVL. Prompted baselines up to Gemma-27B can be run on a single H100; for the Llama-70B prompted baseline we used a machine with 2xH100 (80GB each).
@@
  • Finetuning baselines on the full synthetic dataset: $\sim 3$ hours for finetuning Gemma-12B on a machine with 1xH100 NVL; $\sim 2$ hours for finetuning Llama-8B on a machine with 1xH100; less than two hours for smaller models.
@@ B.1 Generalisation within Synthetic Dataset
  • Length: we generated samples which are very short (~20 words), short (~50 words), medium (~100 words) or long (~200 words).
@@
These were deliberately chosen as factors which could potentially confound our high-stakes target. In Figure 11 we test the generalization of the high-stakes probes across each of these potentially confounding factors, and find that none of them have a significant effect on probe accuracy.
@@ B.2 Probes vs Baselines Across all Datasets
In Figure 12, we see the performance of our probes and all baselines across all test datasets.
For Anthropic, MT, MTS and Toolace, we used a subset of the dataset for development iteration. Aya Redteaming and Mental Health were fully held-out. These are both interesting: for all other datasets, our probes’ performance is roughly in line with the baselines. However our probes significantly outperform all baselines other than Llama-70B on Aya Redteaming, and our probes underperform all baselines (with the exception of Softmax beating Llama-1B finetuned) on Mental Health.
We believe the reason for the performance on Aya Redteaming is that, due to the nature of this dataset, the judgement of high- or low-stakes is particularly fuzzy (see Table 9), and perhaps varies by LLM (which is supported by the fact that both high-scoring classifiers were based on Llama-70B).
@@
The TPR at 1% FPR for our probes is low across all datasets except MT. TPR at 1% FPR is a useful metric since it represents what proportion of positive samples can be caught in deployment with an acceptable cost to users (such as 1% of samples being flagged). As discussed in Section 3, our setting is less sensitive to false positives than a single-layer monitoring setup, thus for internal flagging purposes, operating at 20-40% FPR could be reasonable, where our probes achieve substantially better TPR (see full ROC curves in Figure 4.
Figure 12: Probes achieve AUROC scores comparable to medium-sized prompted and finetuned LLM monitors. We plot the performance of the two probes using activations from Llama-3-70b with the best average performance, and that of the various baselines on the test evaluation datasets. Note: Mental Health and Aya Redteaming datasets did not have dev sets and are thus completely OOD. Probes usually outperform 1B-parameter baselines, while larger baselines (Gemma-27B and Llama-70B) tend to perform best overall. Interestingly, Gemma-12B-finetuned does worse than Llama-8B-finetuned on the two no-dev-set datasets, but performs similarly otherwise—we believe this might be due to limited hyperparameter optimization for Gemma-12B. Error bars represent the standard deviation and are calculated based on training the respective method three times with different random seeds. We did not finetune Gemma-27B and Llama-70B due to technical limitations.
@@ B.3 Comparing alternative methods of combining probes and baselines
  • Selection: After running the probe on all samples, a subset of samples is selected to give to the baseline for further processing. The parameter $k$directly controls the monitoring budget—if we send $k$of samples to the expensive baseline, we incur $k$of the baseline’s computational cost while maintaining 100% coverage through the cheap probe. We consider three strategies: “top”, which selects samples with the highest probe scores, “bottom” which selects samples with lowest probe scores and “mid” which selects samples with scores around the median score.
@@
We tried all these different strategies. In our plots, we only include results for selecting samples with scores around the median and combining them using average, as this way of combination consistently outperformed other strategies. A plot with additional combinations of probe and baseline methods can be found in Figure 13.
The motivation for choosing samples around the median is as follows: given a fixed "expensive monitoring" budget $k$%, we want to select the $k$of samples where the probe is most uncertain. If the calibration curve for the probe is approximately monotonically increasing, then a score far from 0 or 1 should correspond to a lower probability of correctly classifying a sample. Since our probes are not perfectly calibrated (Figure 15), the score with the lowest probability of correctness is not $0.5$. At deployment time, it is reasonable to assume we have access to an approximate distribution of scores (by e.g. keeping track of the last $n$ scores), but it is not reasonable to assume access to the ground truth, so as a reasonable estimate we pick the median score as that which has lowest probability of correctness. As long as the errors of the probe scores are sufficiently uncorrelated with the errors of the baseline scores, combining these probe scores with a baseline score should improve the probability of predicting correctly.
We note that, in a real deployment environment, a cascaded monitoring system would look very different to this experiment. For the purposes of this paper, we chose to focus exclusively on high- and low-stakes, but in reality high-stakes would not be the terminal task, and the purpose of classifying an input as high-stakes would be to flag for further processing. Thus, even though in this experiment we are sending kof samples around the median to a more expensive classifier, that is not something we would recommend in a deployment system.
Figure 13: Results for other combinations of probe and baseline. For many combinations, probe + baseline achieves a better performance than either of the methods alone, which suggests that the methods learn some completementary information. Probe + prompted gemma-1b is the only combination where performance becomes much lower as compared to probe alone.
@@ B.4 Do Probes Outperform Simple Word Statistics?
As can be seen in Figure 14, this simple classifier performs significantly worse on the eval datasets than our Attention probe.
@@ B.5 Calibration of Probes and Baselines
In Figure 15 we see the calibration curve for the three best probe designs on each evaluation dataset. Some datasets show better calibration than others: in particular, Mental Health and MTS have particularly poor calibration. For MTS, this is likely due to the small sample size for each bin. For Mental Health, the poor calibration, taken together with the low AUROC and TPR at 1% FPR relative to baselines in Figure 12, suggests our probes failed to robustly learn the notion of high-stakes in this dataset.
@@
In Figures 16 and 17, we see that calibration is relatively poor for most finetuned and prompted baselines.
@@ B.6 Alternative Strategies for Training on Deployment Samples
In Figure 18 we compare the performance of our probes finetuned on deployment samples (after training on the synthetic dataset) to the performance we get when training a probe directly on synthetic and deployment data combined.
@@ C.1 Details on Synthetic Dataset Generation
We follow the methodology as shown in Figure 19, and described in three steps as follows:
Situation Generation: We first prompt GPT-4o to generate high- and low-stakes situations, varying our prompt along the axes - domains, user roles, and impact of the situations as described in Table 7. These variations ensured diversity in the situations generated. The following is the specific prompt for generating an example situation pair in the environmental domain, with teachers and students as the user roles, and considering the impact of permanent harm.
@@
Sample Generation: For each of the generated situations, we instruct GPT-4o to produce varied and realistic samples representing queries or messages someone might pose to an LLM in those situations. Variations in the samples included altering the tone, length, language, and style. Table 8 summarises the variations and the number of samples corresponding to each of the variation categories. The following is the specific prompt for generating high- and low-stakes sample pairs based on the situation pairs mentioned above.
@@ C.2 Synthetic Dataset Examples
    Label how high-stakes the samples are on scale from 1-10 and how confident this prediction is using GPT-4o, we use the same prompt for filtering our synthetic dataset (see section C.4).
@@
    Create a balanced version of the dataset by removing ambiguous cases (high-stakes score in 4-7 or confidence $<6$)

### v3

@@ A.2 Baselines
For our baselines, we designed four different prompt templates and evaluated the baselines using all four templates on the development split of our evaluation datasets. For each model, we then picked the template that led to the highest average AUROC (see Figure˜10) and evaluated it on the test split of the evaluation datasets. Note that, especially for smaller models, the choice of prompt template can have a great effect on the performance of prompted baselines.
The prompt templates we chose for each model can be found in Table˜4. Below you can find the four prompt templates:
@@
Llama-3.2-1B default
Llama-3.1-8B default
Llama-3.3-70B single-word
Gemma-3-1B prompt-at-end
Gemma-3-12B single-word
Gemma-3-27B single-word
@@
Hyperparameters used for training our finetuned baselines can be found in Figure˜8. We computed a validation loss on the validation split of our synthetic dataset to decide which checkpoint to use. Note that we while we trained for 10-25 epochs for most models, the best epoch (and checkpoint we used for evaluation) was between 2-4 for smaller models (Llama-3.2-1B and Gemma-3-1B), 3-5 for Llama-3.1-8B and 1-2 for Gemma-3-12B. In particular, learning rate scheduling did not help for smaller models. For Llama-3.1-8B, we found gradient clipping to be particularly important for getting good results. Surprisingly, for Gemma-3-12B, a gradient accumulation of 4 led to much better performance on the dev split than using a gradient accumulation of 128, mostly because of the ToolACE dataset (AUROC of around 0.9 compared to AUROC below 0.7).
@@ A.3 Estimating Computational Cost
  • Activations from Llama-3.3-70B for the given conversation (in exactly that conversation format) are already available at no extra cost. This is because we assume the monitored system already uses Llama-3.3-70B as its primary model to generate responses to user queries, and the activations we need are a natural byproduct of this inference process.
    Note that the prompted baseline of Llama-3.3-70B uses a different prompt structure, therefore we assume that this baseline model has to process the whole sequence. (The cost of the prompted Llama-3.3-70B could potentially be reduced by keeping the same conversation format and asking about the stakes label at the end, but it is unclear how much that would affect performance.)
@@
Plugging in the parameters for the specific models, we obtain the compute costs formulas listed in Table˜5.
Model NN $m$ $d$ $w$ Compute Cost Formula
@@
Note that finetuned and prompted baselines use different prompt templates, which explains why the number of FLOPs for the finetuned and prompted baseline of the same model can be different (e.g. see Figure˜13).
@@
Note that we estimate the number of FLOPs for our probes as well, but they are several orders of magnitude smaller and thus not noticeable in probe+baseline combinations in the plot (Figure˜6).
The specific estimates for the relevant probes can be found in Table˜6.
@@
Table 6: Compute cost formulas for different probes on Llama-3.3-70B activations as functions of input sequence length $S$. Note that activations of Llama-3.3-70B have size $d=8192$.
@@ A.4 Hardware Resources
  • Cloud servers from the cloud provider Lambda Labs, Vast.ai and Runpod were used for our experiments. For computing activations, a machine with 2xH100 GPUs (80GB each) and $\sim 400$GB of RAM was used. For training and evaluating probes with pre-computed activations, a machine with 1xA100 (40GB) and $\sim 200$GB of RAM was sufficient. For finetuning baselines up to 8B parameters, a machine with 1xH100 (80GB) was used. Gemma-3-12B was finetuned on a machine with 1xH100 NVL. Prompted baselines up to Gemma-3-27B can be run on a single H100; for the Llama-3.3-70B prompted baseline we used a machine with 2xH100 (80GB each).
@@
  • Finetuning baselines on the full synthetic dataset: $\sim 3$ hours for finetuning Gemma-3-12B on a machine with 1xH100 NVL; $\sim 2$ hours for finetuning Llama-3.1-8B on a machine with 1xH100; less than two hours for smaller models.
@@
A.5 Local LLMs Studied
In this paper, we use several abbreviations to refer to the LLMs we use in our experiments. For the avoidance of confusion, here we provide a table listing the full model names with their corresponding HuggingFace repository URLs.
Abbreviation in Paper Full Model Name HuggingFace URL
Llama-70B / Llama-3-70B llama-3.3-70b-instruct meta-llama/Llama-3.3-70B-Instruct
Llama-8B / Llama-3-8B llama-3.1-8b-instruct meta-llama/Llama-3.1-8B-Instruct
Llama-1B / Llama-3-1B llama-3.2-1b-instruct meta-llama/Llama-3.2-1B-Instruct
Gemma-27B gemma-3-27b-it google/gemma-3-27b-it
Gemma-12B gemma-3-12b-it google/gemma-3-12b-it
Gemma-1B gemma-3-1b-it google/gemma-3-1b-it
Table 7: Mapping of abbreviated model names to full model identifiers used throughout this paper. All models listed are instruction-tuned variants ("-it" suffix for Gemma, "-instruct" suffix for Llama).
Throughout the paper, we omit the instruction-tuning suffix ("-it" for Gemma models, "-instruct" for Llama models) for brevity, and may further abbreviate model names in figures and tables.
@@ B.1 Generalisation within Synthetic Dataset
  • Length: we generated samples which are very short (˜20 words), short (˜50 words), medium (˜100 words) or long (˜200 words).
@@
These were deliberately chosen as factors which could potentially confound our high-stakes target. In Figure˜11 we test the generalization of the high-stakes probes across each of these potentially confounding factors, and find that none of them have a significant effect on probe accuracy.
@@ B.2 Probes vs Baselines Across all Datasets
In Figure˜12, we see the performance of our probes and all baselines across all test datasets.
For Anthropic, MT, MTS and Toolace, we used a subset of the dataset for development iteration. Aya Redteaming and Mental Health were fully held-out. These are both interesting: for all other datasets, our probes’ performance is roughly in line with the baselines. However our probes significantly outperform all baselines other than Llama-3.3-70B on Aya Redteaming, and our probes underperform all baselines (with the exception of Softmax beating Llama-3.2-1B finetuned) on Mental Health.
We believe the reason for the performance on Aya Redteaming is that, due to the nature of this dataset, the judgement of high- or low-stakes is particularly fuzzy (see Table˜11), and perhaps varies by LLM (which is supported by the fact that both high-scoring classifiers were based on Llama-3.3-70B).
@@
The TPR at 1% FPR for our probes is low across all datasets except MT. TPR at 1% FPR is a useful metric since it represents what proportion of positive samples can be caught in deployment with an acceptable cost to users (such as 1% of samples being flagged). As discussed in Section˜3, our setting is less sensitive to false positives than a single-layer monitoring setup, thus for internal flagging purposes, operating at 20-40% FPR could be reasonable, where our probes achieve substantially better TPR (see full ROC curves in Figure˜4).
Figure 12: Probes achieve AUROC scores comparable to medium-sized prompted and finetuned LLM monitors. We plot the performance of the two probes using activations from Llama-3.3-70B with the best average performance, and that of the various baselines on the test evaluation datasets. Note: Mental Health and Aya Redteaming datasets did not have dev sets and are thus completely OOD. Probes usually outperform 1B-parameter baselines, while larger baselines (Gemma-3-27B and Llama-3.3-70B) tend to perform best overall. Interestingly, Gemma-3-12B-finetuned does worse than Llama-3.1-8B-finetuned on the two no-dev-set datasets, but performs similarly otherwise—we believe this might be due to limited hyperparameter optimization for Gemma-3-12B. Error bars represent the standard deviation and are calculated based on training the respective method three times with different random seeds. We did not finetune Gemma-3-27B and Llama-3.3-70B due to technical limitations.
@@ B.3 Comparing alternative methods of combining probes and baselines
  • Selection: After running the probe on all samples, a subset of samples is selected to give to the baseline for further processing. The parameter kk% directly controls the monitoring budget—if we send kk% of samples to the expensive baseline, we incur kk% of the baseline’s computational cost while maintaining 100% coverage through the cheap probe. We consider three strategies: “top”, which selects samples with the highest probe scores, “bottom” which selects samples with lowest probe scores and “mid” which selects samples with scores around the median score.
@@
We tried all these different strategies. In our plots, we only include results for selecting samples with scores around the median and combining them using average, as this way of combination consistently outperformed other strategies. A plot with additional combinations of probe and baseline methods can be found in Figure˜13.
The motivation for choosing samples around the median is as follows: given a fixed "expensive monitoring" budget kk%, we want to select the kk% of samples where the probe is most uncertain. If the calibration curve for the probe is approximately monotonically increasing, then a score far from 0 or 1 should correspond to a lower probability of correctly classifying a sample. Since our probes are not perfectly calibrated (Figure˜15), the score with the lowest probability of correctness is not $0.5$. At deployment time, it is reasonable to assume we have access to an approximate distribution of scores (by e.g. keeping track of the last $n$ scores), but it is not reasonable to assume access to the ground truth, so as a reasonable estimate we pick the median score as that which has lowest probability of correctness. As long as the errors of the probe scores are sufficiently uncorrelated with the errors of the baseline scores, combining these probe scores with a baseline score should improve the probability of predicting correctly.
We note that, in a real deployment environment, a cascaded monitoring system would look very different to this experiment. For the purposes of this paper, we chose to focus exclusively on high- and low-stakes, but in reality high-stakes would not be the terminal task, and the purpose of classifying an input as high-stakes would be to flag for further processing. Thus, even though in this experiment we are sending k% of samples around the median to a more expensive classifier, that is not something we would recommend in a deployment system.
Figure 13: Results for other combinations of probe and baseline. For many combinations, probe + baseline achieves a better performance than either of the methods alone, which suggests that the methods learn some completementary information. Probe + prompted Gemma-3-1B is the only combination where performance becomes much lower as compared to probe alone.
@@ B.4 Do Probes Outperform Simple Word Statistics?
As can be seen in Figure˜14, this simple classifier performs significantly worse on the eval datasets than our Attention probe.
@@ B.5 Calibration of Probes and Baselines
In Figure˜15 we see the calibration curve for the three best probe designs on each evaluation dataset. Some datasets show better calibration than others: in particular, Mental Health and MTS have particularly poor calibration. For MTS, this is likely due to the small sample size for each bin. For Mental Health, the poor calibration, taken together with the low AUROC and TPR at 1% FPR relative to baselines in Figure˜12, suggests our probes failed to robustly learn the notion of high-stakes in this dataset.
@@
In Figures˜16 and 17, we see that calibration is relatively poor for most finetuned and prompted baselines.
@@ B.6 Alternative Strategies for Training on Deployment Samples
In Figure˜18 we compare the performance of our probes finetuned on deployment samples (after training on the synthetic dataset) to the performance we get when training a probe directly on synthetic and deployment data combined.
@@
B.7 Probes on Different Model Sizes
In the main paper, all probes are trained on activations from Llama-3.3-70B (80 layers). To investigate whether probe performance depends on model size, we trained attention probes on two additional models: Llama-3.1-8B (32 layers) and Gemma-3-27B (62 layers). For each model, we selected the probing layer via cross-validation over every 5th layer on the synthetic training dataset, using the same methodology as for Llama-3.3-70B (Section˜A.1).
Results. The Llama-3.1-8B probe achieved a mean AUROC of 0.887 across our six evaluation datasets, while the Gemma-3-27B probe achieved 0.913. For comparison, the Llama-3.3-70B probe achieved 0.916. These results demonstrate that our probing method generalizes across model sizes, with only modest performance degradation on smaller models. Notably, both smaller-model probes still outperform all 1B parameter finetuned and prompted baselines (see Figure˜3).
Performance gaps between probes and their respective base models. An interesting pattern emerges when comparing probe performance to prompted baselines of the same model family. For Llama-3.1-8B, the gap between the probe (0.887) and the prompted baseline (0.877, Figure˜3) is smaller than for Llama-3.3-70B, where the probe (0.916) underperforms the prompted baseline (0.929). This suggests that probes may be particularly valuable for smaller models where prompted baselines are less effective.
Deployment considerations. These results strengthen the practical applicability of our approach. In deployment scenarios where computational constraints require using smaller models for the primary task, probes can still provide effective monitoring. Moreover, the assumption that base model activations are "free" (since they’re generated during normal inference) holds regardless of model size, making probes an attractive monitoring solution across the model scale spectrum.
B.8 Chain-of-Thought Baselines
We evaluated chain-of-thought (CoT) prompting as an alternative baseline, instructing models to reason within scratchpad tags before classification. Table˜8 compares CoT with our standard prompted baselines across various model sizes, including advanced reasoning models.
Table 8: Chain-of-thought prompting performance. Mean AUROC scores across all six test evaluation datasets. CoT substantially improves only Gemma-3-12B. Advanced models (DeepSeek V3, Claude Sonnet 4.0) perform well but do not exceed our optimized Llama-3.3-70B prompted baseline. Invalid response rates shown in parentheses.
Model Prompted Baseline CoT Baseline
Llama-3.2-1B 0.766 0.544 (77.3% invalid)
Gemma-3-1B 0.713 0.597 (35.3% invalid)
Llama-3.1-8B 0.916 0.901
Gemma-3-12B 0.908 0.959
Gemma-3-27B 0.945 0.955
Llama-3.3-70B 0.970 0.928
DeepSeek V3 0324 0.958 (2.8% invalid)
Claude Sonnet 4.0 0.960 (0.1% invalid)
CoT baselines are impractical for real-time monitoring due to: (1) high latency from sequential token generation, and (2) instruction-following failures in smaller models. Our prompted baselines avoid invalid responses by using log-likelihoods (Section 2.2). Interestingly, while advanced models (DeepSeek, Claude) achieve strong performance, they do not exceed our carefully optimized Llama-3.3-70B prompted baseline, likely due to differences in interpreting the inherently subjective concept of “high-stakes” and limited prompt optimization for CoT during our experiments.
@@ C.1 Details on Synthetic Dataset Generation
We follow the methodology as shown in Figure˜19, and described in three steps as follows:
Situation generation: We first prompt GPT-4o to generate high- and low-stakes situations, varying our prompt along the axes - domains, user roles, and impact of the situations as described in Table˜9. These variations ensured diversity in the situations generated. The following is the specific prompt for generating an example situation pair in the environmental domain, with teachers and students as the user roles, and considering the impact of permanent harm.
@@
Sample Generation: For each of the generated situations, we instruct GPT-4o to produce varied and realistic samples representing queries or messages someone might pose to an LLM in those situations. Variations in the samples included altering the tone, length, language, and style. Table˜10 summarises the variations and the number of samples corresponding to each of the variation categories. The following is the specific prompt for generating high- and low-stakes sample pairs based on the situation pairs mentioned above.
@@ C.2 Synthetic Dataset Examples
    Label how high-stakes the samples are on scale from 1-10 and how confident this prediction is using GPT-4o, we use the same prompt for filtering our synthetic dataset (see section˜C.4).
@@
    Create a balanced version of the dataset by removing ambiguous cases (high-stakes score in 4-7 or confidence <6<6)